# Open-Ended 3D Point Cloud Instance Segmentation

### v1

@@ 3.4 Our Approach
To address the above limitations, we propose a method for producing pointwise 3D visual tokens, as illustrated in Fig. 2. First, we generate class-agnostic 2D instance segmentation masks for all views using class-agnostic 2D segmenters such as DETIC detic and SAM sam . Next, we lift 2D masks into 3D masks using Open3DIS nguyen2023open3dis . Simultaneously, these 2D masks, along with their corresponding RGB images, are used to extract 2D visual tokens from an MLLM like OSM yu2023towards . We then lift the resulting 2D visual tokens $\mathbf{F}^{\text{2D}}$ into 3D visual tokens to obtain pointwise 3D visual tokens $\mathbf{F}^{\text{3D}}$. Finally, for each 3D proposal mask, we query the 3D visual tokens associated with this proposal by aggregating the pointwise lifted 3D visual tokens, forming the final tokens $\mathbf{f}^{\text{3D}}$ for input to the LLM.
This approach takes into account the depth and geometric structure of 3D objects via lifting, resulting in more robust visual tokens and a unified, densely-featured point cloud that can be queried instantly at test time. Subsequently, we will focus on our pointwise visual tokens lifting and aggregation.
First, the correspondence of a 3D point $\mathbf{p}_{i}(x,y,z)\in\mathbf{P}$ with its 2D projection $(u,v)$ in view $t$ is:
@@
where $d_{i,t}$ is the projected 2D depth value of point $i$ to frame $t$.
For each view $t$, we extract the 2D visual tokens $\mathbf{F}^{k,\text{2D}}_{t}\in\mathbb{R}^{E\times C}$, where $E,C$ are the number of visual tokens and feature dimensions, for each 2D mask $k$ from the MaskQ-Former module of OSM yu2023towards . Furthermore, we denote $\lambda^{k}_{t,i}=\{0,1\}$ as the visibility value indicating a point $i$ is visible in mask $k$ of view $t$, $\lambda^{k}_{t}=\{0,1\}^{N}$. We set the visibility value $\lambda_{t,i}^{k}$ of any points whose pixel projections fall outside the $k$-th 2D mask’s boundaries or the disparity between projected depth $d$ and the collected depth $\mathbf{D}$ exceeds a defined depth threshold $\tau_{\text{depth}}$, or $|d_{i,t}-\mathbf{D}_{t}[\lfloor u_{i}\rfloor,\lfloor v_{i}\rfloor]|>\tau_{depth}$, to 00.
We accumulate 3D visual tokens $\mathbf{F}^{\text{3D}}\in\mathbb{R}^{N\times E\times C}$, from every 2D mask $k$ and compute for the frequency $\mathbf{r}^{\text{3D}}\in\mathbb{N}^{N}$ of every view as follows:
$\mathbf{F}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t}*\mathbf{F}^{k,\text{2D}}_{t},\quad\quad\quad\quad\quad\quad\mathbf{r}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t},$ (4)
where * is the outer product. The normalized pointwise 3D visual tokens are then obtained as follows:
@@ 4 Experimental Results
Datasets: We conducted experiments to assess the performance of the baselines and our proposed method using two common 3DIS datasets: ScanNet200 rozenberszki2022language and ScanNet++ yeshwanthliu2023scannetpp . The ScanNet200 dataset builds on the original ScanNet dai2017scannet by expanding its semantic categories from 20 to 200. Its instance segmentation benchmark includes 1,201 training scenes and 312 validation scenes with 198 object categories, significantly enriching the vocabulary and enhancing its capability for detailed 3D semantic and instance segmentation. The ScanNet++ dataset was recently introduced, featuring up to 1,659 semantic categories, with 360 training scenes and 50 validation scenes. Given the large number of classes, we follow the standard 3DIS evaluation protocol on ScanNet++ and evaluate only the most common 100 object categories. This dataset offers a much denser 3D point cloud scene representation, making it the most challenging dataset for 3D understanding.
@@
Implementation details: Following Open3DIS nguyen2023open3dis , we generate class-agnostic 3D proposals from ISBNet ngo2023isbnet pretrained on ScanNet200 or by lifting 2D masks to 3D using Detic detic . For RAM++ zhang2023recognize , we employ the Swin-L model, trained on a 14-million image dataset with an image size of 384px and a tagging threshold of 0.68. For LLM, we leverage Vicuna-7B vicuna2023 , fine-tuned for open-ended 2D Instance Segmentation yu2023towards .
Table 1: Comparative results on the ScanNet200 dataset. Shaded text indicates a reference method, not a direct comparison. ‘-’ indicates results are not provided. The best results are in bold.
@@
Mask3D Schult23mask3d 26.9 36.2 41.4 39.8 21.7 17.9 -
Fully-sup 3DIS ISBNet ngo2023isbnet 24.5 32.7 37.6 38.6 20.5 12.5 -
OpenMask3D takmaz2023openmask3d 15.4 19.9 23.1 17.1 14.1 14.9 -
OV-3DIS Open3DIS nguyen2023open3dis 23.7 29.4 32.8 27.8 21.2 21.8 -
Large-Vocab 21K DETIC classes (Ours) 8.5 11.7 13.1 9.9 7.2 8.3 52.2
Image-Tagging RAM++ zhang2023recognize (Ours) 10.7 14.3 16.0 11.6 11.0 9.3 60.5
Maskwise OSM yu2023towards (Ours) 14.4 19.8 23.9 18.9 13.5 10.2 65.0
Pointwise Ours 16.0 22.0 24.7 20.0 14.3 13.2 65.2
Table 2: Comparative results on the ScanNet++ dataset. Shaded text indicates a reference method, not a direct comparison. ‘-’ indicates results are not provided. The best results are in bold.
Setting Method AP AP50 AP25 AR RC50 RC25 OE
PointGroup jiang2020pointgroup 8.9 14.6 38.9 - - - -
Fully-sup 3DIS SoftGroup vu2022softgroup 16.7 29.7 21.0 - - - -
OpenMask3D takmaz2023openmask3d 2.0 2.7 3.4 4.6 8.3 12.4 -
OV-3DIS Open3DIS (known classes) nguyen2023open3dis 13.1 20.8 24.6 22.1 33.9 39.1 -
Large-Vocab 21K DETIC classes (Ours) 7.3 11.9 15.2 13.3 20.3 23.6 42.0
Image-Tagging RAM++ zhang2023recognize (Ours) 9.1 15.5 19.1 16.0 24.8 28.7 48.5
Maskwise OSM yu2023towards (Ours) 16.3 24.8 29.0 22.2 32.0 36.0 51.2
Pointwise Ours 18.4 29.4 33.6 23.3 35.2 39.3 53.9
4.1 Comparison with Baselines
We compare our approach to the proposed baselines using the OE-3DIS setting on the ScanNet200 (Tab. 1) and ScanNet++ (Tab. 2) datasets. For reference, we also present results from OV-3DIS methods, including OpenMask3D takmaz2023openmask3d and Open3DIS nguyen2023open3dis , as well as fully-supervised methods like ISBNet ngo2023isbnet and Mask3D Schult23mask3d for ScanNet200; and PointGroup jiang2020pointgroup and SoftGroup vu2022softgroup for ScanNet++.
For Scannet200: We obtain class-agnostic 3D proposals from two sources: 3D masks from a 3DIS network such as ISBNet ngo2023isbnet , and 2D Lift 3D masks from Open3DIS nguyen2023open3dis . Our proposed approach outperforms other baselines in both AP and OE scores. Furthermore, the performance progression of the baselines, in the specified order, clearly justifies the motivation behind each baseline compared to its predecessor, as discussed in Sec. 3.3. Interestingly, our approach also surpasses OpenMask3D takmaz2023openmask3d (16.0 vs. 15.4 in AP), even though OpenMask3D utilizes provided class names. This indicates that, in some cases, we can achieve OV-3DIS without relying on provided class names.
For ScanNet++: Due to the extensive scale and vast array of classes in ScanNet++, the performance of 3D mask results from ISBNet is inadequate. Consequently, we solely rely on the utilization of 2D Lift 3D masks from Open3DIS nguyen2023open3dis . We notice a consistent trend akin to the results observed in ScanNet200. Particularly noteworthy is the significant outperformance of our approach compared to OV-3DIS methods or even fully-supervised 3DIS, as evidenced by higher AP scores. This underscores the promising application of OE-3DIS in navigating complex 3D scene structures.
Qualitative comparison: In Fig. 3, both our proposed baselines and approach effectively assign class names to 3D proposals in OE-3DIS settings. The first column of ScanNet200 qualitative results shows our architectures accurately predicting all 3D proposals. The second column highlights differences between our approach and the baselines. Our method precisely identifies the ‘painting’ object as “wall painting”, while other baselines predict less accurate object labels. In the final column, our 3D mask proposals offer more detail than ground truth, enabling predictions of classes not in ScanNet++’s vocabulary, such as “photocopier”.
Figure 3: Qualitative examples are provided for ScanNet200 dai2017scannet (first two columns) and ScanNet++ yeshwanthliu2023scannetpp (last two columns). Both our baselines and approach yield notably good results, particularly in terms of accurately identifying class names even though they do not match exactly the GT classes.
@@ 4.1 Comparison with Baselines
Max 6.3 8.4 10.2 8.5 4.7 5.6 56.3
Random 13.2 18.9 21.6 17.7 12.3 10.0 60.4
Mean 14.5 20.1 22.6 18.2 13.1 11.2 63.1
Weighted Average 16.0 22.0 24.7 20.0 14.3 13.2 65.2
@@
3D masks DETIC 11.7 16.0 18.3 16.3 10.3 8.1 56.4
2D Lift 3D masks DETIC 11.7 18.6 23.5 11.3 12.1 11.8 54.3
3D + 2D Lift 3D masks DETIC 16.0 22.0 24.7 20.0 14.3 13.2 65.2
3D + 2D Lift 3D masks SAM 15.4 20.5 22.9 19.2 14.8 11.7 64.8
@@ 4.2 Ablation Study
To study many design choices of our pointwise approach, we intensively carry out ablation study on the ScanNet200 dai2017scannet dataset.
Study on different point feature aggregation techniques (Eq. 6) is shown in Tab. 3. We evaluate four techniques for combining point features from a given 3D mask: max, random (randomly selecting one point), mean, and weighted average (our proposed operation). The weighted average technique achieved the highest performance with an AP score of 16.0, outperforming the other methods. Consequently, we utilized the weighted average technique in all our experiments.
Study on different types 3D proposals. As reported in Tab. 4, combining 3D proposals from both 3D masks and 2D Lift 3D masks yielded the most favorable results. This outcome validates our choice of 3D proposals for the ScanNet200 dataset. Additionally, the 2D Lift 3D proposals obtained from DETIC demonstrated slightly superior performance compared to those from SAM. This observation diverges from the methodology adopted in Open3DIS nguyen2023open3dis , where provided class names are utilized to employ Grounded-SAM ren2024grounded instead of SAM sam , as in our OE-3DIS scenario.
Study on different text encoders for evaluation metrics is described in Tab. 5. We assess three text decoders: BERT devlin2018bert , CLIP radford2021learning , and Sentence Transformer sentence-transformer . Among them, the Sentence Transformer embeddings attain the highest AP scores, whereas the CLIP text embeddings achieve the highest OE score. The OE score indicates that the text embeddings from CLIP exhibit a higher degree of similarity to each other, contrasting with those from the Sentence Transformer, which tend to describe complex sentences rather than solely focusing on a single class name. Hence, we lean towards selecting the Sentence Transformer as our recommended text encoder.
@@
Table 5: Study on different text encoders used in evaluation metrics
[…]
BERT devlin2018bert 13.5 19.0 21.2 17.6 11.6 11.0 70.6
CLIP radford2021learning 15.9 22.0 24.5 20.3 13.7 13.3 71.6
Sentence Transformer sentence-transformer 16.0 22.0 24.7 20.0 14.3 13.2 65.2
Table 6: Study on different text prompts used to query the LLM.
Text Prompt AP OE
“What is in the segmentation mask? Assistant:” 16.0 65.2
“Can you describe what is in the segmentation mask region? Assistant:” 15.3 64.3
“What can you see in the segmentation mask region? Assistant:” 15.9 65.0
“Could you use a few words to describe what is in the segmentation mask region? Assistant:” 15.2 64.0
“What is this segmentation mask? Assistant:” 15.8 65.0
@@ 5 Discussion and Conclusion
Limitations: Though our method enhances open-ended 3D scene understanding, it has notable drawbacks. Firstly, it heavily relies on 2D visual tokens from the pretrained OSM, which are trained on instance segmentation datasets with limited classes. This hampers the model’s ability to recognize a wide range of classes in an open-world environment. Secondly, the performance of OE-3DIS is contingent upon class-agnostic 3D instance segmentation (3DIS) methods, which in turn rely on the quality of 3D representation and 2D-3D mapping, such as camera calibrations and depth images.
Conclusion: We have introduced Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS), which generates 3D masks and object class names without predefined labels during testing. We have explored baselines using OV-3DIS methods and MLLMs, and introduced a pointwise training-free approach leveraging OSM. We have also proposed the OE score, tailored for OE-3DIS. Experiments on ScanNet200 and ScanNet++ show our approach’s superior performance, notably outperforming Open3DIS (SOTA on OV-3DIS) on ScanNet++ without ground-truth class names.

### v2

@@ 3.4 Our Approach
To address the above limitations, we propose a method for producing pointwise 3D visual tokens, as illustrated in Fig. 3. First, we generate class-agnostic 2D instance segmentation masks for all views using class-agnostic 2D segmenters such as DETIC [63] and SAM [31]. Next, we lift 2D masks into 3D masks using Open3DIS [39]. Simultaneously, these 2D masks, along with their corresponding RGB images, are used to extract 2D visual tokens from an MLLM like OSM [59]. We then lift the resulting 2D visual tokens $\mathbf{F}^{\text{2D}}$ into 3D visual tokens to obtain pointwise 3D visual tokens $\mathbf{F}^{\text{3D}}$. Finally, for each 3D proposal mask, we query the 3D visual tokens associated with this proposal by aggregating the pointwise lifted 3D visual tokens, forming the final tokens $\mathbf{f}^{\text{3D}}$ for input to the LLM. This approach takes into account the depth and geometric structure of 3D objects via lifting, resulting in more robust visual tokens and a unified, densely-featured point cloud that can be queried instantly at test time. Subsequently, we will focus on our pointwise visual tokens lifting and aggregation.
Concretely, first, the correspondence of a 3D point $\mathbf{p}_{i}(x,y,z)\in\mathbf{P}$ with its 2D projection $(u,v)$ in view $t$ is:
@@
where $d_{i,t}$ is the projected depth of point $i$ to frame $t$.
Next, for each view $t$, we extract the 2D visual tokens $\mathbf{F}^{k,\text{2D}}_{t}\in\mathbb{R}^{E\times C}$, where $E,C$ are the number of visual tokens and feature dimensions, for each 2D mask $k$ from the MaskQ-Former module of OSM [59]. Furthermore, we denote $\lambda^{k}_{t,i}=\{0,1\}$ as the visibility value indicating a point $i$ is visible in mask $k$ of view $t$, $\lambda^{k}_{t}=\{0,1\}^{N}$. We set the visibility value $\lambda_{t,i}^{k}$ of any points whose pixel projections fall outside the $k$-th 2D mask’s boundaries or the disparity between projected depth $d$ and the collected depth $\mathbf{D}$ exceeds a defined depth threshold $\tau_{\text{depth}}$, or $|d_{i,t}-\mathbf{D}_{t}[\lfloor u_{i}\rfloor,\lfloor v_{i}\rfloor]|>\tau_{depth}$, to 0.
Then, we accumulate 3D visual tokens $\mathbf{F}^{\text{3D}}\in\mathbb{R}^{N\times E\times C}$, from every 2D mask $k$ and compute for the frequency $\mathbf{r}^{\text{3D}}\in\mathbb{N}^{N}$ of every view as follows:
$\mathbf{F}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t}*\mathbf{F}^{k,\text{2D}}_{t},\quad\quad\mathbf{r}^{\text{3D}}=\sum_{t,k}\lambda^{k}_{t},$ (2)
where * is the outer product operation. The normalized pointwise 3D visual tokens are then obtained as follows:
@@ 4 Experimental Results
Datasets: We conducted experiments to assess the performance of the baselines and our proposed method using two common 3DIS datasets: ScanNet200 [49] and ScanNet++ [56]. The ScanNet200 dataset builds on the original ScanNet [17] by expanding its semantic categories from 20 to 200. Its instance segmentation benchmark includes 1,201 training scenes and 312 validation scenes with 198 object categories, significantly enriching the vocabulary and enhancing its capability for detailed 3D semantic and instance segmentation. The ScanNet++ dataset was recently introduced, featuring up to 1,659 semantic categories, with 360 training scenes and 50 validation scenes. Given the large number of classes, we follow the standard 3DIS evaluation protocol on ScanNet++ and evaluate only the most common 100 object categories. This dataset offers a much denser 3D point cloud scene representation, making it the most challenging dataset for 3D understanding.
@@
Implementation details: Following Open3DIS [39], we generate class-agnostic 3D proposals from ISBNet [37] pretrained on ScanNet200 or by lifting 2D masks to 3D using Detic [63]. For RAM++ [29], we employ the Swin-L model, trained on a 14-million image dataset with an image size of 384px and a tagging threshold of 0.68. For LLM, we leverage Vicuna-7B [9], fine-tuned for open-ended 2D Instance Segmentation [59].
Setting Method ScanNet200 ScanNet++
AP AP50 AP25 $_{\text{h}} AP$$_{\text{c}}$ $_{\text{t}} AP$ AP50 AP25 AR RC50 RC25
Mask3D [50] 26.9 36.2 41.4 39.8 21.7 17.9 8.9 14.6 38.9 - - -
Fully-sup 3DIS ISBNet [37] 24.5 32.7 37.6 38.6 20.5 12.5 16.7 29.7 21.0 - - -
OpenMask3D [51] 15.4 19.9 23.1 17.1 14.1 14.9 2.0 2.7 3.4 4.6 8.3 12.4
OV-3DIS Open3DIS [39] 23.7 29.4 32.8 27.8 21.2 21.8 13.1 20.8 24.6 22.1 33.9 39.1
Large-Vocab 21K DETIC classes (Ours) 8.5 11.7 13.1 9.9 7.2 8.3 7.3 11.9 15.2 13.3 20.3 23.6
Image-Tagging RAM++ [29] (Ours) 10.7 14.3 16.0 11.6 11.0 9.3 9.1 15.5 19.1 16.0 24.8 28.7
Maskwise OSM [59] (Ours) 14.4 19.8 23.9 18.9 13.5 10.2 16.3 24.8 29.0 22.2 32.0 36.0
Pointwise Ours 16.0 22.0 24.7 20.0 14.3 13.2 18.4 29.4 33.6 23.3 35.2 39.3
Table 1: Comparative results on the ScanNet200 and ScanNet++ datasets. Shaded text indicates a reference method, not a direct comparison. ‘-’ indicates results are not provided. The best results are in bold.
4.1 Comparison with Baselines
We compare our approach to the proposed baselines using the OE-3DIS setting on the ScanNet200 and ScanNet++ datasets in Tab. 1. For reference, we also present results from OV-3DIS methods, including OpenMask3D [51] and Open3DIS [39], as well as fully-supervised methods like ISBNet [37] and Mask3D [50] for ScanNet200; and PointGroup [30] and SoftGroup [52] for ScanNet++.
For Scannet200: We obtain class-agnostic 3D proposals from two sources: 3D masks from a 3DIS network such as ISBNet [37], and 2D Lift 3D masks from Open3DIS [39]. Our proposed approach outperforms other baselines in the AP score. Furthermore, the performance progression of the baselines, in the specified order, clearly justifies the motivation behind each baseline compared to its predecessor, as discussed in Sec. 3.3. Interestingly, our approach also surpasses OpenMask3D [51] (16.0 vs. 15.4 in AP), even though OpenMask3D utilizes provided class names. This indicates that, in some cases, we can achieve OV-3DIS without relying on provided class names.
For ScanNet++: Due to the extensive scale and vast array of classes in ScanNet++, the performance of 3D mask results from ISBNet is inadequate. Consequently, we solely rely on the utilization of 2D Lift 3D masks from Open3DIS [39]. We notice a consistent trend akin to the results observed in ScanNet200. Particularly noteworthy is the significant outperformance of our approach compared to OV-3DIS methods or even fully supervised 3DIS, as evidenced by higher AP scores. This underscores the promising application of OE-3DIS in navigating complex 3D scene.
Qualitative comparison: In Fig. 4, both our proposed approach and baseline methods effectively assign class names to 3D proposals under OE-3DIS settings. The first column of qualitative results on ScanNet200 demonstrates that our architectures correctly predict class labels for all 3D proposals. The second column highlights notable differences, with our method accurately identifying the ‘painting’ object as a ‘wall painting’, whereas the baselines produce less precise labels. In the final column, our detailed 3D mask proposals surpass the granularity of the ground-truth annotations, enabling accurate recognition of classes absent from ScanNet++’s vocabulary, such as ‘photocopier’. This illustrates how OE-3DIS allows the model to comprehensively understand a 3D scene without restricting to fixed categories.
@@ 4.1 Comparison with Baselines
L2 Norm (Open3DIS) 5.7 8.2 9.8 6.8 4.9 5.2
Memory Fusion (OVIR-3D) 8.4 11.6 14.4 7.2 7.0 11.6
Max 6.3 8.4 10.2 8.5 4.7 5.6
Random 13.2 18.9 21.6 17.7 12.3 10.0
Mean 14.5 20.1 22.6 18.2 13.1 11.2
Weighted Average (Ours) 16.0 22.0 24.7 20.0 14.3 13.2
@@
3D masks DETIC 11.7 16.0 16.3 10.3 8.1
2D Lift 3D masks DETIC 11.7 18.6 11.3 12.1 11.8
3D + 2D Lift 3D masks DETIC 16.0 22.0 20.0 14.3 13.2
3D + 2D Lift 3D masks SAM 15.4 20.5 19.2 14.8 11.7
@@
BERT [20] 13.5 19.0 21.2 17.6 11.6 11.0
CLIP [43] 15.9 22.0 24.5 20.3 13.7 13.3
Sentence Transformer [44] 16.0 22.0 24.7 20.0 14.3 13.2
Table 4: Study on different text encoders in evaluation metrics
Modules # Params(M) FLOPs (G) Time (s)
Pw. CLIP extract (CLIP) 427.94 191.11 97
QFormer 185.66 2.74 250
LLM (Vicuna 7B) 6,738.42 850.01 50
RAM++ 329.49 104.43 40
Generate 3D proposals N/A 357.99 152
Pw. Visual Token Lift (PVTL) N/A 4.915 314
Table 5: Expected latency, FLOPs, and runtime of each module for a 3D scene. Pw. denotes ‘Pointwise’. Generating 3D proposals from 2D and Pw. visual token lifting are non-parametric modules
Figure 4: Qualitative examples are provided for ScanNet200 [17] (first two columns) and ScanNet++ [56] (last two columns). Both our baselines and approach yield notably good results, particularly in terms of accurately identifying class names even though they do not match exactly the GT classes.
Method AP Total Time (s) RAM++ Generate 3D proposals CLIP QFormer PVTL LLM
Open3DIS N/A 249
Large Vocab 8.5 249
Image Tagging 10.7 289
Maskwise 14.4 452
Pointwise 16.0 766
Table 6: Expected runtime for our proposed baselines and methods with the Open3DIS for a sample 3D scene
Text Prompt AP
“What is in the segmentation mask? Assistant:” 16.0
“Can you describe what is in the segmentation mask region? Assistant:” 15.3
“What can you see in the segmentation mask region? Assistant:” 15.9
“Could you use a few words to describe what is in the segmentation mask? Assistant:” 15.2
“What is this segmentation mask? Assistant:” 15.8
Table 7: Study on different text prompts used to query the LLM.
@@ 4.2 Ablation Study
To study many design choices of our pointwise approach, we intensively carry out ablation study on the ScanNet200 [17] dataset.
Study on different point feature aggregation techniques (Eq. 4) is shown in Tab. 2. We evaluate four techniques for combining point features from a given 3D mask: L2-norm, max, random (randomly selecting one point), mean, and weighted average (our proposed operation). The weighted average technique achieved the highest performance with an AP score of 16.0, outperforming the other methods. Consequently, we utilized the weighted average technique in all our experiments.
Study on different types 3D proposals. As reported in Tab. 3, combining 3D proposals from both 3D masks and 2D Lift 3D masks yielded the most favorable results. This outcome validates our choice of 3D proposals for the ScanNet200 dataset. Additionally, the 2D Lift 3D proposals obtained from DETIC demonstrated slightly superior performance compared to those from SAM. This observation diverges from the methodology adopted in Open3DIS [39], where provided class names are utilized to employ Grounded-SAM [45] instead of SAM [31], as in our OE-3DIS scenario.
Study on different text encoders for evaluation metrics is described in Tab. 4.We evaluate three text decoders: BERT [20], CLIP [43], and Sentence Transformer [44]. Among these, Sentence Transformer embeddings achieve the highest AP scores, outperforming CLIP and BERT, which yield comparatively lower scores. We also observe that CLIP embeddings exhibit greater similarity among themselves compared to those from BERT. Moreover, Sentence Transformer embeddings are better suited for capturing complex sentence-level descriptions rather than focusing solely on individual class names. Therefore, we recommend Sentence Transformer as our preferred text encoder.
Latency analysis of our baselines and proposed method is presented in Tab. 5 (excluding the runtime for generating 2D proposals). Additionally, Tab. 6 compares the total runtime across our baselines, proposed methods, and the state-of-the-art 3DIS approach, Open3DIS. As shown, our method prioritizes accuracy at the cost of increased runtime, meaning higher accuracy comes with slower execution.
@@ 5 Discussion and Conclusion
Limitations: The motivation behind our proposed Open-ended 3D Instance Segmentation (OE-3DIS) arises from limitations in current Open-Vocabulary 3D Point Cloud Instance Segmentation (OV-3DIS) methods, which still depend on a predefined set of class names during testing. This constraint is impractical in scenarios lacking prior knowledge of class names, such as a robot navigating unfamiliar environments. Although our method advances open-ended 3D scene understanding, it still faces notable limitations. Firstly, our approach heavily depends on 2D visual tokens extracted from a pretrained OSM, which itself is trained on instance segmentation datasets containing only a limited number of classes. Consequently, this restricts the model’s capability to recognize an extensive range of classes in truly open-world contexts. Secondly, the performance of OE-3DIS relies significantly on class-agnostic 3DIS methods, whose effectiveness in turn depends on the accuracy of 3D representations and the quality of 2D-to-3D mapping, including factors like camera calibration and depth image quality.
Conclusion: We have introduced Open-Ended 3D Point Cloud Instance Segmentation (OE-3DIS), which generates 3D masks and object class names without predefined labels during testing. We have explored baselines using OV-3DIS methods and MLLMs, and introduced a pointwise training-free approach leveraging OSM. Experiments on ScanNet200 and ScanNet++ show our approach’s superior performance, notably outperforming Open3DIS (SOTA on OV-3DIS) on ScanNet++ without ground-truth class names.